OpenStudy (anonymous):

We want to construct a box whose base length is 3 times the base width. The material used to build the top and bottom cost $10/ft2 and the material used to build the sides cost $6/ft2. If the box must have a volume of 50ft3 determine the dimensions that will minimize the cost to

OpenStudy (anonymous):

AB = 3w*w area,base TB=2*3*w^2 Total base area SA= 2*3w*x+w*x*2 side areas all 4 50=3w*2*x eq for volume price = 10/AB price = 6/SA for side areas hold on i'll work it out :)